# Differentially Private Adaptation of Diffusion Models via Noisy Aggregated Embeddings

### v3

@@ C.4 Ablation
Appendix D Copyright Protection
Modern generative models typically produce outputs via randomized sampling. Leveraging this inherent randomness, Vyas, Kakade, and Barak (2023) introduced Near Access-Freeness (NAF) as a metric to quantify the similarity between a model’s output and copyrighted content. The key idea is to compare the output distribution of a potentially infringing model to that of a safe model – one trained without access to the target content.
Formally, let $\mathsf{safe}$ be a mapping from a data point $x\in\mathcal{C}$ (where $\mathcal{C}$ is the collection of copyrighted samples) to a generative model $\mathsf{safe}(x)\in\mathcal{W}$ that is trained without using $x$. A canonical example is the leave-one-out-safe model, trained on the full dataset excluding $x$. Since $\mathsf{safe}(x)$ does not have access to $x$, the probability that it generates content resembling $x$ is exponentially small. Any such resemblance is considered fortuitous. Formally, the NAF criterion is defined as follows:
Definition 2 (Near Access-Freeness (Vyas, Kakade, and Barak 2023)).
Let $\mathcal{C}$ be a set of copyrighted samples and $\mathcal{W}$ a set of generative models. Given a mapping $\mathsf{safe}:\mathcal{C}\rightarrow\mathcal{W}$ and a divergence measure $\Delta$, we say a model $w$ is $k_{y}$-near access-free (or $k_{y}$-NAF) on prompt $y\in\mathsf{Y}$ if for every $x\in\mathcal{C}$,
$\Delta\left(p(\cdot|y)\,\|\,\mathsf{safe}_{x}(\cdot|y)\right)\leq k_{y}.$
D.1 Connection to Differential Privacy
NAF is closely related to concepts in Differential Privacy (DP) (Elkin-Koren et al. 2023). Depending on the divergence measure $\Delta$, NAF resembles different DP variants – for example, $\varepsilon$-DP when $\Delta=\Delta_{\mathrm{max}}$ (Dwork et al. 2006), and $(1,\varepsilon)$-Rényi DP when $\Delta=\Delta_{\mathrm{KL}}$.
Translating DP to generative models yields the following definition:
Definition 3 (Differentially Private Generation (DPG)).
Let $S$ and $S^{\prime}$ be neighboring datasets. Denote by $P_{S}(\cdot|y)$ the distribution over outputs generated by a model trained on, or adapted from, $S$ with algorithm $\mathcal{A}$, where randomness includes both training and generation stages. The generation is said to satisfy $\varepsilon$-Differentially Private Generation ($\varepsilon$-DPG) if for every $y\in\mathcal{Y}$,
Here, neighboring datasets differ by a single data point (or privacy unit). If the training process is $\varepsilon$-DP, then the outputs naturally satisfy $\varepsilon$-DPG via the data processing inequality. One benefit of DPG is the flexibility to add noise during generation rather than training, potentially improving the utility-privacy tradeoff.
However, there are notable distinctions. $\varepsilon$-DP offers protection under arbitrary post-processing and multiple outputs, whereas $\varepsilon$-DPG only guarantees privacy for single outputs. Also, under DP, the trained model can be released, but under DPG, only the outputs are safe to share.
Elkin-Koren et al. (2023) highlight further differences: NAF is one-sided—comparing a model to a fixed safe reference—whereas DPG is symmetric. This asymmetry in NAF can enable better utility. Additionally, NAF allows more flexibility in choosing the safe model, which can be exploited in algorithm design.
Given these conceptual overlaps, both DP-SGD based training and our proposed private adaptation method DPAgg-TI satisfies $\varepsilon$-DP, so they naturally satisfy $\varepsilon$-NAF with the leave-one-out safe model.
We emphasize that this guarantee is meaningful only within the formal framework of NAF. It does not imply broader legal immunity or empirical indistinguishability from the original content. However, within this framework, satisfying $\varepsilon$-NAF allows us to argue that any close resemblance between outputs and private training data is no more likely than would be expected from a model that never had access to that data. This theoretical grounding supports the privacy and safety claims of our adaptation method.
Importantly, the goal of DPAgg-TI is to adapt to the style of a private image set—not its precise content. This distinction matters: style transfer is widely considered to fall under the doctrine of fair use, particularly in artistic and creative contexts. As discussed in Elkin-Koren et al. (2023) and further elaborated in legal analysis such as Carlini et al. (2023), generating new content in the style of a work, without reproducing its substantive elements, is generally not considered copyright infringement. Therefore, the use of DPAgg-TI to learn and reproduce stylistic attributes does not contradict the spirit or intent of the NAF framework. Instead, it offers a promising direction for responsibly fine-tuning generative models on private or copyrighted sources while respecting both privacy and intellectual property boundaries.
Appendix E Computational Cost Comparisons
@@
Direct comparisons of computational cost across methods are inherently challenging due to differing training paradigms, optimization procedures, and parameter settings. Nonetheless, to provide a concrete sense of scale, we report representative computational costs for each method based on experiments conducted using a Stable Diffusion v1.5 model on a single NVIDIA A100 GPU. Below we summarize both training and inference overheads (the number of steps are optimized for each setup):
@@
Table 4: Training cost comparison across methods. Overhead from DP-SGD is relatively modest due to the low-dimensional embedding being optimized. N/A for SG means nothing is trained aside from the base model.
@@
Table 5: Inference cost comparison across methods.
Appendix F Additional Style Transfer and Ablation Results

### v4

@@ A-D Ablation
Appendix B Computational Cost Comparisons
@@
Direct comparisons of computational cost across methods are challenging due to differing training paradigms (per-image optimization vs. dataset-level training), optimization procedures, and privacy accounting. Nonetheless, to provide a concrete sense of scale, we report representative costs measured using Stable Diffusion v1.5 on a single NVIDIA A100 GPU (Tables IV and V). For each method, we tuned the number of optimization steps to reach its best utility under the target privacy budget.
Sequential vs. batched execution
The per-image runtime reported in Table IV corresponds to a sequential implementation that optimizes each textual-inversion embedding independently. In practice, we can optimize multiple embeddings jointly by batching several images at once (and optionally subsampling the private set per update), which reduces the effective wall-clock cost and avoids the naive linear scaling implied by “minutes per image $\times\;n$.”
TABLE IV: Training cost comparison across methods. Overhead from DP-SGD is relatively modest due to the low-dimensional embedding being optimized. N/A for SG means nothing is trained aside from the base model.
@@
TABLE V: Inference cost comparison across methods.
@@
When DP-SGD can be faster
DP-SGD amortizes computation across the dataset and can be faster wall-clock-wise for larger $n$ and/or moderate privacy budgets. In contrast, our method is designed for low-data personalization with strong per-record guarantees, and it offers a practical advantage in dynamic settings: it supports incremental updates to the private set (e.g. adding or removing images) without retraining a large set of model parameters, whereas DP-SGD-style training typically requires rerunning optimization to reflect such changes. We view these approaches as complementary, targeting different operating regimes.
Appendix C Additional Style Transfer and Ablation Results